# Impact of in-situ nuclear networks and atomic opacities on neutron star merger ejecta dynamics, nucleosynthesis, and kilonovae

### v1

@@ 2.4 Thermalization
The thermalization factor $f_{j}^{\gamma}(t)$ for prompt $\gamma$-rays is calculated, as in Hotokezaka & Nakar (2019); Combi & Siegel (2023), starting from the detailed composition of the ejecta and the same effective opacity tables from the NIST-XCOM catalogue (Berger et al., 2010) used in Barnes et al. (2016). The thermalization efficiency associated with the $\gamma$ particles emitted by nuclei of the $j$-th isotope through radioactive decays is given by
@@
where $Z$ runs over all possible atomic numbers, $X_{Z}=\sum_{s:Z_{s}=Z}X_{s}$, with $s$ running through all the isotopes considered in the NN, and $\kappa_{\text{eff},Z}^{\gamma}(E_{k,j})$ is the effective opacity for a stopping material made only of isotopes with atomic number $Z$. The latter is included in the NIST-XCOM catalogue. Since this database only contains information up to $Z=100$, we fix $\kappa_{\text{eff},Z>100}^{\gamma}(E_{k,j})=\kappa_{0}$ and use Eq. (15) to rewrite Eq. (14) as
@@
For the remaining nuclear power, associated with the recoil energy of daughter nuclei, fission fragments, (soft) $X$-rays, proton and neutron emissions, and delayed electron (e.g. Auger) emission, we assume a constant and uniform thermalization factor $f_{\text{th}}^{\text{oth}}$. The heavy charged particles are expected to thermalize efficiently at all times, leaving only to $X$-rays, injected nucleons and delayed electrons with a non-trivial behavior. In our simulations, we choose, as a standard value, $f_{\text{th}}^{\text{oth}}=0.8>f_{\text{th}}^{\text{{\tt kNEC}}}$, as we are now explicitly removing the neutrino contributions. The total thermalized heating rate is finally given by
@@ 2.5 Opacities
For models based on the full tracking of the matter composition, we rely on the line-binned opacity calculations presented in Fontes et al. (2020, 2023), and available online via the NIST-LANL Opacity Database (Ralchenko et al., 2025). In these publications, the population distribution of the ionization states is computed for a range of temperatures and densities using the Saha equation. Detailed ab-initio atomic-structure calculations are then employed to compute the opacity as a function of discretized frequency intervals. No assumptions are needed about the dynamical expansion. In this work, we use the resulting opacities for a series of selected atoms, and use them to effectively define a composition-averaged opacity of the ejecta. We discretize the frequency domain $\nu\in[3\times 10^{9},\penalty 10000\ 3.6\times 10^{19}]$ Hz in $n_{\textrm{gr}}=300$ groups. We use a piecewise logarithmic frequency grid with coarse spacing at low ($\nu\lesssim 2.42\times 10^{12}$ Hz) and high ($\nu\gtrsim 7.25\times 10^{15}$ Hz) energies, and higher resolution in the central energy region to better resolve spectral features. We then calculate the group opacity from the line-binned results as prescribed in (Fontes et al., 2020). The database includes data for $10^{-20}<\rho[\text{g}/\text{cm}^{3}]<10^{-4}$ and $116\lesssim T[\text{K}]\lesssim 5.8\times 10^{4}$. For densities or temperatures above the limits, we use the analytical model from Just et al. (2022). Close to the upper boundaries of the table, we use an activation function to smoothly switch between the two models. For densities or temperatures below the limits, we use the opacity data at the lower edge of the table. We use Nd and U data as representatives for all the lanthanides or actinides, respectively. All the remaining elements are assigned an opacity $\kappa_{\text{oth}}=0.2\penalty 10000\ k_{T}$ cm2g-1, with $k_{T}=1$ if $T>2000$ K and $k_{T}=(T[\textrm{K}]/2000)^{5}$ for $T<2000$ K, consistent with the no-lanthanides case of the Just et al. (2022) approximation. To get the composition-averaged opacity $\bar{\kappa}_{\nu}$ of Eq. (3), we first estimate the opacity $\kappa_{\nu,i}(\rho,T)$ of each representative element by interpolating its tabulated values in the $\rho-T$ plane. Then, for each frequency group, we average the contributions of the representative elements, weighting each element by the total mass fractions of the species it represents. The composition-averaged opacity is given by
@@ 2.6 Jet Energy Deposition
If a GRB is launched from the polar regions of the remnant, it releases additional energy into the ejecta it traverses. kNECnn can mimic this effect by adding an extra energy term to the right-hand side of Eq. (1) for the innermost shells in the polar regions of the ejecta. We use the “thermal bomb” model already implemented in SNEC (Morozova et al., 2015), and define two parametrized prescriptions for the energy released by the GRB.
In the “simple jet” model, we assume a jet with an opening angle $\theta_{j}$ to release a uniform isotropic energy $E_{\text{iso}}(\theta)=E_{0}$ in the traversed regions. The total energy released by the jet is given by $E_{j}=(\Delta\Omega_{j}/4\pi)E_{0}$, where $\Delta\Omega_{j}=2\pi(1-\cos\theta_{j})$ is the solid angle covered by the jet (or two times this value if one assumes reflection with respect to the equatorial plane).
@@ 2.7 Initial profiles
Our simulations are initialized with initial ejecta profiles extracted from ab-initio NR simulations performed with the THC code and including a microphysical equation of state (EOS), M1 neutrino transport, and magnetic-field induced turbulence (Radice et al., 2018; Perego et al., 2019; Bernuzzi, 2020; Radice et al., 2022; Zappa et al., 2023; Bernuzzi et al., 2025). The main parameters of the considered binaries are listed in Table 1. All models have already been presented in Bernuzzi et al. (2025); Gutierrez et al. (2025). For details on the EOS used, see Hempel & Schaffner-Bielich (2010); Typel et al. (2010); Bombaci & Logoteta (2018); Logoteta et al. (2021). Both selected models BLh_1.43 and DD2_1.67 describe asymmetric BNSM. The chirp mass of the BLh_1.43 configuration is compatible with GW170817, and the mass ratio lies at the upper bound of the constraints inferred with low spin priors Abbott et al. (2019). The DD2_1.67 represent a comparable but more extreme mass-ratio scenario. The neutron stars composing the binary systems are all non-rotating. The central remnant is a massive neutron star for the duration of the simulation in both models.
@@
Figure 2: Thermodynamic profile of the ejecta at the beginning of the kNECnn simulation for the BLh_1.43 binary. From left to right and top to bottom, the quadrants show the initial density, temperature, entropy and electron fraction.
Table 1: Main parameters of the ab-initio NR simulations from which we generate the initial profiles for the kNECnn runs discussed in this paper. From left to right, the name of the simulation, the EOS employed, the mass ratio $q$, the gravitational masses of the two non-rotating neutron stars, $M_{1}$ and $M_{2}$, the final time $t_{f}$ of the simulation, the time we choose to distinguish between dynamical and wind ejecta $t_{\text{dyn}}$, and the mass of these two components, $M_{\text{dyn}}$ and $M_{\text{wind}}$. The times are reported here as times after merger. The remnants do not collapse to a black hole before the simulations end.
@@ 3 Results
Table 2: List of kNECnn simulations analyzed in this paper. From left to right, the name of the simulation and of the NR-extracted initial ejecta profile (see Table 1), the NSE initialization method employed (BK6 and BK8 stand for the backtracking procedure starting at $T=6,8$ GK, CNSE for the cold-NSE method; Apr2 simulations are performed without in-situ NN), the thermalization scheme used (DT08 and ST05 stand for the detailed thermalization scheme described in Sec. 2.4 with $f_{\text{th}}^{\text{oth}}=0.8$ or the simple thermalization from Wu et al. (2022) with $f_{\text{th}}^{\text{{\tt kNEC}}}=0.5$). The next columns state wether a frequency-dependent (and with how many frequency groups) or a gray opacity is used, which model is employed (LANL stands for the atomic calculations, J[22] for the analytical model from Just et al. (2022), and W[22] for the opacity from Wu et al. (2022)), and which atomic opacities are read from the NIST-LANL database (see Sec. 2.5). The last three columns report the peak of the bolometric luminosity, the time of this peak, and the final mass fraction of lanthanides produced.
@@ 3.1 Effects of dynamics on nucleosynthesis
Figure 3: Mass-weighted nucleosynthesis yields at $t\simeq 5\times 10^{4}$ s for the BLh (left column) and DD2 (right column) ejecta profiles. Top panels: Final yields for the simulations performed with in-situ NN (blue) or using analytical fits for the nuclear power and post-processing the nucleosynthesis either on the original thermodynamic trajectories (orange), or on analytical density evolutions prescribed from a grid of initial thermodynamic conditions, with a standard (green) or double (red) resolution on the $Y_{e},\tau,s$ grid. The black dots represent the Solar $r$-process abundances (Prantzos et al., 2020). All the abundances are scaled to produce a unitary total abundance of the elements with $170\leq A\leq 200$. The histogram shows, for the in-situ NN run, the global cumulative mass fractions of the first (blue), second (orange) and third (purple) $r$-process peaks, and the rare-earths (brown). Bottom panel: relative differences against the simulation with in-situ NN. The upwards [downwards] triangles indicate discrepancies greater than two orders of magnitude [smaller than $5\times 10^{-4}$]. The horizontal dashed lines highlight the $1\%$ and factors of 1 and 10 discrepancies.
@@
The pp_traj analysis reproduces most of the final abundances within a $\sim 20\%$ error. This improvement comes from the more accurate representation of the density (and temperature) evolution along each trajectory. If (opacity and) thermalization effects are neglected, and if reliable analytical fits for the nuclear powers are employed, a sufficiently large sample of tracers, well resolved in time up to $\sim 1$ s, can therefore yield reliable nucleosynthesis predictions (if non-radial motion is neglected). The remaining discrepancies originate from corrections to the thermodynamic evolution of the fluid elements that come exclusively from the improved nuclear-power calculation provided by the in-situ NN (see next section for details).
@@
In Fig. 4, we compare the time evolution of selected species obtained from the runs with in-situ NN to those from the post-processing tracers procedure. We focus on neutrons and protons as diagnostic isotopes, 56Ni (half-life $\tau\simeq 6.077$ days), 92Sr (with $\tau\simeq 2.611$ hr), 128Sb (with $\tau\simeq 9.05$ hr) and 131I (with $\tau\simeq 8.02$ days) for their potential contributions to $\gamma$-ray emission (Korobkin et al., 2019; Bernuzzi et al., 2025; Jacobi et al., 2025), 60Fe, 129I, 244Pu and 247Cm for their relevance in galactic chemical evolution studies (Côté et al., 2021; Wallner et al., 2021; Davis, 2022; Chiesa et al., 2024), and the cumulative mass fractions of lanthanides and actinides to monitor strong $r$-process nucleosynthesis. At $t\sim 1$ s, the system departs from $(n,\gamma)\leftrightarrow(\gamma,n)$ equilibrium. The neutron-to-seed ratio consequently drops, triggering neutron freeze-out. The remaining free neutrons undergo $\beta$-decay on the $t\sim 10$ minutes timescale. 56Ni and lanthanides are already present at $t=0$. 56Ni is further synthesized on the $t\sim 10$ ms timescale, while lanthanides are produced for $t\gtrsim 100$ ms together with actinides. The isotopes 128Sb, 129I, and 131I are produced at significant abundances on a timescale of $t\sim 1$ hour, with 128Sb already decaying at $t\sim 10$ hours.
@@ 3.2 Effects of nucleosynthesis on dynamics and light curves
The differences in the predicted evolution of the photosphere temperature and the nuclear power released in optically thin regions, however, have a qualitative impact on the light curves. In Fig. 6, we compare kilonova predictions computed with and without in-situ NN for the BLh and DD2 ejecta profiles. The light curves are brighter, peak later, and last longer in the DD2 case, which produces more than double the amount of ejected mass, as expected from the Arnett’s law (Arnett, 1982). Both simulations show a blue peak at $t\sim 1$ hour, when the photosphere enters regions previously heated up by the decay of freshly synthesized $r$-process isotopes (Bernuzzi et al., 2025). The late-time red kilonova is powered both by the ongoing decays of these neutron-rich isotopes, and by the 56Ni $\to^{56}$Co $\to^{56}$Fe decay chain (Jacobi et al., 2025).
@@
Calculating the nuclear power on the fly generally shifts the emission peak to later times and lower frequencies. The lower luminosity at $t\sim 1$ hour is caused by the colder temperature predicted by the online NN for the photosphere, which most contributes to the light curves on this timescale. We show this effect in Fig. 7, which displays the heating rate and temperature profiles for the BLh simulation at $t\sim 1$ hour. At this time, nuclear contributions from the optically thin layers of the ejecta provide only a minor correction to the luminosity, and differences between the in-situ and ex-situ simulations arise only in low-activity regions. Similar considerations hold for the DD2 profile.
Figure 7: Thermalized heating rate and temperature profiles of the BLh ejecta at $t\sim 1$ hour. The top panels show the results obtained with in-situ NN, and the bottom panels show the relative differences with respect to the run using analytical nuclear power fits.

### v2

@@ 2.4 Thermalization
The thermalization factor $f_{j}^{\gamma}(t)$ for prompt $\gamma$-rays is calculated, as in Hotokezaka & Nakar (2019); Combi & Siegel (2023), starting from the detailed composition of the ejecta and the same effective opacity tables from the NIST-XCOM catalog (Berger et al., 2010) used in Barnes et al. (2016). The thermalization efficiency associated with the $\gamma$ particles emitted by nuclei of the $j$-th isotope through radioactive decays is given by
@@
where $Z$ runs over all possible atomic numbers, $X_{Z}=\sum_{s:Z_{s}=Z}X_{s}$, with $s$ running through all the isotopes considered in the NN, and $\kappa_{\text{eff},Z}^{\gamma}(E_{k,j})$ is the effective opacity for a stopping material made only of isotopes with atomic number $Z$. The latter is included in the NIST-XCOM catalog. Since this database only contains information up to $Z=100$, we fix $\kappa_{\text{eff},Z>100}^{\gamma}(E_{k,j})=\kappa_{0}$ and use Eq. (15) to rewrite Eq. (14) as
@@
For the remaining nuclear power, associated with the recoil energy of daughter nuclei, fission fragments, (soft) $X$-rays, proton and neutron emissions, and delayed electron (e.g. Auger) emission, we assume a constant and uniform thermalization factor $f_{\text{th}}^{\text{oth}}$. The heavy charged particles are expected to thermalize efficiently at all times, leaving only to $X$-rays, injected nucleons and delayed electrons with a nontrivial behavior. In our simulations, we choose, as a standard value, $f_{\text{th}}^{\text{oth}}=0.8>f_{\text{th}}^{\text{{\tt kNEC}}}$, as we are now explicitly removing the neutrino contributions. The total thermalized heating rate is finally given by
@@ 2.5 Opacities
For models based on the full tracking of the matter composition, we rely on the line-binned opacity calculations presented in Fontes et al. (2020, 2023), and available online via the NIST-LANL Opacity Database (Ralchenko et al., 2025). In these publications, the population distribution of the ionization states is computed for a range of temperatures and densities using the Saha equation. Detailed ab-initio atomic-structure calculations are then employed to compute the opacity as a function of discretized frequency intervals. No assumptions are needed about the dynamical expansion. kNECnn implements data for Cr, Se, Br, Zr, Pd, Te, Ce, Nd, Sm and U. In this work, we focus on the effect of using atomic data for lanthanides and actinides, and use only the resulting opacities for Nd and U to effectively define a composition-averaged opacity of the ejecta. We discretize the frequency domain $\nu\in[3\times 10^{9},\penalty 10000\ 3.6\times 10^{19}]$ Hz in $n_{\textrm{gr}}=300$ groups. We use a piecewise logarithmic frequency grid with coarse spacing at low ($\nu\lesssim 2.42\times 10^{12}$ Hz) and high ($\nu\gtrsim 7.25\times 10^{15}$ Hz) energies, and higher resolution in the central energy region to better resolve spectral features. We then calculate the group opacity from the line-binned results as prescribed in (Fontes et al., 2020). The database includes data for $10^{-20}<\rho[\text{g}/\text{cm}^{3}]<10^{-4}$ and $116\lesssim T[\text{K}]\lesssim 5.8\times 10^{4}$. For densities or temperatures above the limits, we use the analytical model from Just et al. (2022). Close to the upper boundaries of the table, we use an activation function to smoothly switch between the two models. For densities or temperatures below the limits, we use the opacity data at the lower edge of the table. We use Nd and U data as representatives for all the lanthanides or actinides, respectively. All the remaining elements are assigned an opacity $\kappa_{\text{oth}}=0.2\penalty 10000\ k_{T}$ cm2g-1, with $k_{T}=1$ if $T>2000$ K and $k_{T}=(T[\textrm{K}]/2000)^{5}$ for $T<2000$ K, consistent with the no-lanthanides case of the Just et al. (2022) approximation. To get the composition-averaged opacity $\bar{\kappa}_{\nu}$ of Eq. (3), we first estimate the opacity $\kappa_{\nu,i}(\rho,T)$ of each representative element by interpolating its tabulated values in the $\rho-T$ plane. Then, for each frequency group, we average the contributions of the representative elements, weighting each element by the total mass fractions of the species it represents. The composition-averaged opacity is given by
@@ 2.6 Jet energy deposition
If a GRB is launched from the polar regions of the remnant, it releases additional energy into the ejecta it traverses. kNECnn can mimic this effect by adding an extra energy term to the right-hand side of Eq. (1) for the innermost shells in the polar regions of the ejecta. We use the “thermal bomb” model already implemented in SNEC (Morozova et al., 2015). The energy of the bomb is released during a chosen time interval and is exponentially attenuated both in time and with distance from the inner boundary. We define two parametrized prescriptions for the energy released by the GRB.
In the “simple jet” model, we assume a jet with an opening angle $\theta_{j}$ to release a uniform (in $\theta$) isotropic energy $E_{\text{iso}}(\theta)=E_{0}$ in the traversed regions. The total energy released by the jet is given by $E_{j}=(\Delta\Omega_{j}/4\pi)E_{0}$, where $\Delta\Omega_{j}=2\pi(1-\cos\theta_{j})$ is the solid angle covered by the jet (or two times this value if one assumes reflection with respect to the equatorial plane).
@@ 2.7 Initial profiles
Our simulations are initialized with initial ejecta profiles extracted from ab-initio NR simulations performed with the THC code and including a microphysical EOS, M1 neutrino transport, and magnetic-field induced turbulence (Radice et al., 2018; Perego et al., 2019; Bernuzzi, 2020; Radice et al., 2022; Zappa et al., 2023; Bernuzzi et al., 2025). The main parameters of the considered binaries are listed in Table 3. All models have already been presented in Bernuzzi et al. (2025); Gutierrez et al. (2025). For details on the EOS used, see Hempel & Schaffner-Bielich (2010); Typel et al. (2010); Bombaci & Logoteta (2018); Logoteta et al. (2021). Both selected models BLh_1.43 and DD2_1.67 describe asymmetric BNSM. The chirp mass of the BLh_1.43 configuration is compatible with GW170817, and the mass ratio lies at the upper bound of the constraints inferred with low spin priors Abbott et al. (2019). The DD2_1.67 represent a comparable but more extreme mass-ratio scenario. The neutron stars composing the binary systems are all nonrotating. The central remnant is a massive neutron star for the duration of the simulation in both models.
@@
In constructing the Lagrangian ejecta profiles, we neglect non-radial components of the ejecta velocity. Moreover, owing to the ray-by-ray nature of the simulations, possible fallback at later times is largely suppressed, since outer shells can only fallback if the inner shells are also infalling. We do not expect these assumptions to introduce strong qualitative corrections to the results discussed in this paper, and defer a quantitative assessment of their impact on the ejecta evolution and final observables to future work.
@@
Figure 2: Thermodynamic profile of the ejecta at the beginning of the kNECnn simulation for the BLh_1.43 binary. From left to right, the quadrants show the initial density, temperature, entropy, radial velocity, and electron fraction.
Table 1: Main parameters of the ab-initio NR simulations from which we generate the initial profiles for the kNECnn runs discussed in this paper. [^3] From left to right, the name of the simulation, the EOS employed, the mass ratio $q$, the gravitational masses of the two non-rotating neutron stars, $M_{1}$ and $M_{2}$, the final time $t_{f}$ of the simulation, the time we choose to distinguish between dynamical and wind ejecta $t_{\text{dyn}}$, and the mass of these two components, $M_{\text{dyn}}$ and $M_{\text{wind}}$. The times are reported here as times after merger. The remnants do not collapse to a black hole before the simulations end.
@@ 3 Results
Table 2: List of kNECnn simulations analyzed in this paper.  [^4] From left to right, the name of the simulation and of the NR-extracted initial ejecta profile (see Table 3), the NSE initialization method employed (BK6 and BK8 stand for the backtracking procedure starting at $T=6,8$ GK, CNSE for the cold-NSE method; Apr2 simulations are performed without in-situ NN), the thermalization scheme used (DT08 and ST05 stand for the detailed thermalization scheme described in Sec. 2.4 with $f_{\text{th}}^{\text{oth}}=0.8$ or the simple thermalization from Wu et al. (2022) with $f_{\text{th}}^{\text{{\tt kNEC}}}=0.5$). The next columns state wether a frequency-dependent (and with how many frequency groups) or a gray opacity is used, which model is employed (LANL stands for the atomic calculations, J[22] for the analytical model from Just et al. (2022), and W[22] for the opacity from Wu et al. (2022)), and which atomic opacities are read from the NIST-LANL database (see Sec. 2.5). The last three columns report the peak of the bolometric luminosity, the time of this peak, and the final mass fraction of lanthanides produced.
@@ 3.1 Effects of dynamics on nucleosynthesis
Figure 3: Mass-weighted nucleosynthesis yields at $t\simeq 5\times 10^{4}$ s for the BLh (left column) and DD2 (right column) ejecta profiles. Top panels: Final yields for the simulations performed with in-situ NN (blue) or using analytical fits for the nuclear power and post-processing the nucleosynthesis either on the original thermodynamic trajectories (orange), or on analytical density evolutions prescribed from a grid of initial thermodynamic conditions, with a standard (green) or double (red) resolution on the $Y_{e},\tau,s$ grid. The black dots represent the Solar $r$-process abundances (Prantzos et al., 2020). All the abundances are scaled to produce a unitary total abundance of the elements with $170\leq A\leq 200$. The histogram shows, for the in-situ NN run, the global cumulative mass fractions of the first (blue), second (orange) and third (purple) $r$-process peaks, and the rare-earths (brown). Bottom panel: relative differences against the simulation with in-situ NN. The upward [downward] triangles indicate discrepancies greater than two orders of magnitude [smaller than $5\times 10^{-4}$]. The horizontal dashed lines highlight the $1\%$ and factors of 1 and 10 discrepancies.
@@
The pp_traj analysis reproduces most of the final abundances within a $\sim 20\%$ error. This improvement comes from the more accurate representation of the density (and temperature) evolution along each trajectory. If (opacity and) thermalization effects are neglected, and if reliable analytical fits for the nuclear powers are employed, a sufficiently large sample of tracers, well resolved in time up to $\sim 1$ s, can therefore yield reliable nucleosynthesis predictions (if non-radial motion is neglected). The remaining discrepancies originate from corrections to the thermodynamic evolution of the fluid elements that come exclusively from the improved nuclear-power calculation provided by the in-situ NN (see next section for details). The effect on the final mass fraction of lanthanides is of the order of $10\%$, as reported in Table 4.
@@
In Fig. 4, we compare the time evolution of selected species obtained from the runs with in-situ NN to those from the post-processing tracers procedure. We focus on neutrons and protons as diagnostic isotopes, 56Ni (half-life $\tau\simeq 6.077$ days), 92Sr (with $\tau\simeq 2.611$ hr), 128Sb (with $\tau\simeq 9.05$ hr) and 131I (with $\tau\simeq 8.02$ days) for their potential contributions to $\gamma$-ray emission (Korobkin et al., 2019; Bernuzzi et al., 2025; Jacobi et al., 2026), 60Fe, 129I, 244Pu and 247Cm for their relevance in galactic chemical evolution studies (Côté et al., 2021; Wallner et al., 2021; Davis, 2022; Chiesa et al., 2024), and the cumulative mass fractions of lanthanides and actinides to monitor strong $r$-process nucleosynthesis. At $t\sim 1$ s, the system departs from $(n,\gamma)\leftrightarrow(\gamma,n)$ equilibrium. The neutron-to-seed ratio consequently drops, triggering neutron freeze-out. The remaining free neutrons undergo $\beta$-decay on the $t\sim 10$ minutes timescale. 56Ni and lanthanides are already present at $t=0$. 56Ni is further synthesized on the $t\sim 10$ ms timescale, while lanthanides are produced for $t\gtrsim 100$ ms together with actinides. The isotopes 128Sb, 129I, and 131I are produced at significant abundances on a timescale of $t\sim 1$ hour, with 128Sb already decaying at $t\sim 10$ hours.
@@ 3.2 Effects of nucleosynthesis on dynamics and light curves
The differences in the predicted evolution of the photosphere temperature and the nuclear power released in optically thin regions, however, have a qualitative impact on the light curves. In Fig. 6, we compare kilonova predictions computed with and without in-situ NN for the BLh and DD2 ejecta profiles. The light curves are brighter, peak later, and last longer in the DD2 case, which produces more than double the amount of ejected mass, as expected from the Arnett’s law (Arnett, 1982). Both simulations show a blue peak at $t\sim 1$ hour, when the photosphere enters regions previously heated up by the decay of freshly synthesized $r$-process isotopes (Bernuzzi et al., 2025). The late-time red kilonova is powered both by the ongoing decays of these neutron-rich isotopes, and by the 56Ni $\to^{56}$Co $\to^{56}$Fe decay chain (Jacobi et al., 2026).
@@
Calculating the nuclear power on the fly generally shifts the emission peak to later times and lower frequencies (see also Table 4). The lower luminosity at $t\sim 1$ hour is caused by the colder temperature predicted by the online NN for the photosphere, which most contributes to the light curves on this timescale. We show this effect in Fig. 7, which displays the heating rate and temperature profiles for the BLh simulation at $t\sim 1$ hour. At this time, nuclear contributions from the optically thin layers of the ejecta provide only a minor correction to the luminosity, and differences between the in-situ and ex-situ simulations arise only in low-activity regions. Similar considerations hold for the DD2 profile.
Figure 7: Thermalized heating rate and temperature profiles of the BLh ejecta at $t\sim 1$ hour. The top panels show the results obtained with in-situ NN, and the bottom panels show the relative differences with respect to the run using analytical nuclear power fits. The dashed black and the dotted gray lines represent the position of the photosphere in the runs with in-situ NN and fits, respectively.